UNIFIED ANALYSIS OF 14 DATA SETS INTERCONNECTING 12 ELECTRONIC STATES OF As_{2}

\underline{CHRISTIAN O. UIBEL}^{a}, Fachbereich 9–Physikalische Chemie, Bergische Universität–Gesamthochschule Wuppertal, 42097 Wuppertal, Germany; ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

The entire emission spectrum of gas phase As_{2} below 32,000 cm^{−1} excited by microwave or high voltage discharges has been measured using a high resolution Fourier-transform spectrometer (Bruker IFS 120HR) with very sensitive Ge or InSb detectors. Its analysis yielded data on fourteen electronic band systems among the twelve electronic states: X 0_{g}^{+}(^{1}Σ), A′ 0_{u}^{−}(^{1}Σ_{u}^{−}), D 1_{g}(^{1}Π_{g}), a 0_{u}^{+}(^{3}Σ_{u}^{−}), a 1_{u}(^{3}Σ_{u}^{−}), c 1_{u}(^{3}Σ_{u}^{+}), c 0_{u}^{−}(^{3}Σ_{u}^{+}), e 1_{u}(^{3}Δ_{u}), e 2_{u}(^{3}Δ_{u}), f 0_{g}^{+}(^{3}Π_{g}), f 0_{g}^{−}(^{3}Π_{g}) and f 1_{g}(^{3}Π_{g}). Rather than analyze these band systems in a series of independent two-state analyses, we have performed a comprehensive combined analysis of all data sets simultaneously, using program DSParFit.^{b} Differences between the present results and those obtained from independent two-state analyses, will be discussed.

a Present address: Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada

b R.J. Le Roy, University of Waterloo Chemical Physics Research Report CP-646 (2000); see also http://leroy.uwaterloo.ca.